Question 9 (7 points) In an English course, students had to turn in a final paper. The mean grade earned by the class was 78.3 with a standard deviation of 3.8 points. The scores are approximately bell- shaped. Approximately 68% of all final paper scores were between two values A and what is the value of A? Write only a number as your answer. Round to one decimal place. Your Answer: Answer Question 10 (7 points) In a sample of Starbucks customers it was found that the number of individual items bought per month at Starbucks approximately bell-shaped. Approximately 95% of the time, the number of monthly items purchased was between two values A and B. What is the value of 8? Write only a number as your answer was 15 with a standard d Your Answer: Answer
Answer #1

Please give a "thumbs up" for the answer , in case you're satisfied with the answer.

9.

Mean = 78.3
Stdev = 3.8

A and B have 68% of all final paper scores. 68% is 1 deviation from mean.

So, A and B are 78.3-3.8 and 78.3+3.8 respectively.

A = 78-3.8 = 74.2
B = 78+3.8 = 81.8